40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jun 10 20:11:46 2022, yuta, Update, ASC, Yarm ASS re-tuning in progress Screenshot_2022-06-10_20-10-52.png
    Reply  Mon Jun 13 20:26:09 2022, yuta, Update, ASC, Yarm ASS re-tuning in progress -part 2- Screenshot_2022-06-13_20-47-12.pngScreenshot_2022-06-13_20-44-43.png
       Reply  Tue Jun 14 20:57:15 2022, Anchal, Update, ASC, Yarm ASS working now 
Message ID: 16911     Entry time: Mon Jun 13 20:26:09 2022     In reply to: 16909     Reply to this: 16915
Author: yuta 
Type: Update 
Category: ASC 
Subject: Yarm ASS re-tuning in progress -part 2- 

[Anchal, Yuta]

We are still in the progress of re-commissioning Yarm ASS.
Today, we tried to adjust output matrix by measuring the sensing matrix at DC.
Turning on yaw loops kind of works, but pitch does not. It seems like there is too much coupling in pitch to yaw.
We might need to adjust the coil output matrix of ITMY and ETMY to go further, and/or try measuring the sensing matrix including pitch - yaw coupling.

What we did:
 - Confirmed that turning on TT1 and TT2 loops (max-transmission loops) work fine. When we intentionally misalign TT1/2, the ASS loops correct it. So, we moved on to measure the sensing matrix of A2L paths, instead of using theoretical matrix caluclated from cavity geometry we used last week (40m/16909).
 - Instead of +/-1's, we put +/-2's in the ITMY coil output matrix to balance the actuation between ETMY and ITMY to take into account that ITMY is now using only two coils for actuating pitch and yaw (40m/16899).
 - Measured the change in C1:ASS-YARM_(E|I)TM_(PIT|YAW)_L_DEMOD_I_OUT16 error signals when offset was added to C1:SUS-(E|I)TMY_ASC(PIT|YAW)_OFFSET. We assumed pitch-yaw coupling is small enough here. Below was the result.

                            ETM PIT error  ITM PIT error
ETM PIT OFFSET of +100cnts: -3.0cnts       -2.99cnts
ITM PIT OFFSET of +100cnts
: -11.94cnts      -5.38cnts

                            ETM PIT error  ITM PIT error
ETM YAW OFFSET of +100cnts:
-3.42cnts      -16.93cnts
ITM YAW OFFSET of +10 cnts: +1.41cnts      +0.543cnts


 - Inverted the matrix to get A2L part of C1:ASS-YARM_OUT_MTRX. Attachment #1 is the current configuration so far.
 - With this, we could close all yaw loops when pitch loops were not on. But vise versa didn't work.
 - Anyway, we aligned the IFO by centering the beams on test masses by our eyes and centered all the oplevs (Attachment #2).

Next:
 - Do coil balancing to reduce pitch-yaw coupling
 - Measure sensing matrix also for pitch-yaw coupling
 - Xarm ASS is also not working now. We need to do similar steps also for Xarm

Attachment 1: Screenshot_2022-06-13_20-47-12.png  132 kB  Uploaded Mon Jun 13 21:47:36 2022  | Hide | Hide all
Screenshot_2022-06-13_20-47-12.png
Attachment 2: Screenshot_2022-06-13_20-44-43.png  242 kB  Uploaded Mon Jun 13 21:48:21 2022  | Hide | Hide all
Screenshot_2022-06-13_20-44-43.png
ELOG V3.1.3-